Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A string of 5 different colored light bulbs is wired in

Expert replies
by BTGmoderatorDC » Tue Aug 27, 2019 5:53 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

A string of 5 different colored light bulbs is wired in such a way that if any two-consecutive light bulb fails, then entire string fails.
If for each individual light bulb, the probability of not failing during time period T is 0.85, what is the probability that the string of light bulbs will fail during the time period T?

Options
a) 0.0225
b) 0.09
c) 0.225
d) 0.25
e) 0.5

OA B

Source: e-GMAT
Join the discussion
Source: — Problem Solving |

by swerve » Thu Aug 29, 2019 2:14 pm
We have to find the probability of fail, on the other hand, we have given that if two consecutive bulbs fail then only the complete string will fail.
So, given probability of not fail \(= 0.85\)

So
probability of fail \(= 0.15\)

In string of 5 in how many ways 2 consecutive will fail \(= 4\) ways (12,23,34,45)

probability of fail \(= 0.15\cdot 0.15 \cdot 4\)
\(= 0.09 \Rightarrow\) __B__
Join the discussion

by Ian Stewart » Fri Aug 30, 2019 7:40 am
BTGmoderatorDC wrote:A string of 5 different colored light bulbs is wired in such a way that if any two-consecutive light bulb fails, then entire string fails.
If for each individual light bulb, the probability of not failing during time period T is 0.85, what is the probability that the string of light bulbs will fail during the time period T?

Options
a) 0.0225
b) 0.09
c) 0.225
d) 0.25
e) 0.5

OA B

Source: e-GMAT
They do not have the correct answer among the answer choices (the correct answer is roughly 0.0794), and this question is way too complicated for the GMAT. I googled the question, and looked at the official explanation posted here. There they essentially just add the probabilities of four cases: bulbs 1+2 fail, or bulbs 2+3 fail, or bulbs 3+4 fail, or bulbs 4+5 fail. So they arrive at the answer (4)(3/20)^2 = 0.09. That is wrong, because the cases overlap: when bulbs 1 and 2 fail, it is sometimes true that bulb 3 also fails. If you just add the probability that 1+2 fail and the probability that 2+3 fail, you're counting the situation when bulbs 1+ 2+3 all fail twice instead of once. You simply can't add probabilities for different cases when cases overlap, unless you account for that overlap.

It's easy to see why their method is wrong just by imagining that each bulb fails 100% of the time. Using the method in the "official explanation", you'd find that the string of bulbs fails 400% of the time, which clearly makes no sense.

To solve the problem properly, avoiding any overlap among cases, you have to break the question down into non-overlapping situations, so you can correctly add the probabilities of each case. This is way too hard for the GMAT, but one way to do it is as follows:

case 1: all five bulbs fail, --the probability of this case is (0.15)^5
case 2: exactly four bulbs fail, which can happen in five ways, so the probability of this case is 5*(0.15)^4 * (0,85)
case 3: exactly three bulbs fail, and we do not have the sequence Fails, Works, Fails, Works, Fails (the only sequence with three failures where you don't have two consecutive failures) -- this can happen in 10 - 1 = 9 ways, so the probability of this case is 9 * (0.85)^2 * (0.15)^3
case 4: exactly two bulbs fail, and those two bulbs are consecutive, which can happen in 4 ways, so the probability of this case is 4 * (0.85)^3 (0.15)^2

Summing those probabilities with a good calculator, you get an answer slightly less than 0.08 (assuming I entered everything correctly). But there's no chance you'd ever need to answer a question like this on the GMAT.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion

by Scott@TargetTestPrep » Thu Sep 05, 2019 5:05 am
BTGmoderatorDC wrote:A string of 5 different colored light bulbs is wired in such a way that if any two-consecutive light bulb fails, then entire string fails.
If for each individual light bulb, the probability of not failing during time period T is 0.85, what is the probability that the string of light bulbs will fail during the time period T?

Options
a) 0.0225
b) 0.09
c) 0.225
d) 0.25
e) 0.5
Since the probability of not failing is 0.85, the probability for failure is 1 - 0.85 = 0.15 for any given light bulb. Thus, if we take any two of these light bulbs, the probability that they both fail is (0.15)^2 = 9/400. Notice that in a string of 5 light bulbs, there are 4 pairs of light bulbs that are consecutive. Thus, the probability that a consecutive pair of light bulbs (and hence the whole system of light bulbs) fails is 4*9/400 = 9/100 = 0.09.

Answer: B

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion